Warner is hiding a 1500 g of goldbar with a temperature of 30 ̊C inside a

tank full of water with a temperature of 97 ̊C. If the mass of water is 30 kg,

calculate the final temperature of the gold bar. ( Specific Heat Capacity of

gold is 129 J/kg ̊C, water is 4200 J/kg ̊C)


1
Expert's answer
2021-01-21T16:14:27-0500
"m_wc_w(T_w-T)=m_gc_g(T-T_g)\\
(30)(4200)(97-T)=(129)(1.5)(T-30)\\T=96.9\degree\ C"